OpenStudy (anonymous):

Solve x2 - 8x - 20 = 0

OpenStudy (mertsj):

Could you factor it?

myininaya (myininaya):

I wonder what 2(-10) is and 2+(-10) is

jhonyy9 (jhonyy9):

x_1,_2=(8+/- sqrt(64+80))/2=(8+/- sqrt144)/2=(8+/-12)/2 = = -4/2 =-2 and 20/2 =10